# Handling Label Noise via Instance-Level Difficulty Modeling and Dynamic Optimization

### v3

@@ Abstract
Recent studies indicate that deep neural networks degrade in generalization performance under label noise. Existing methods focus on data selection or label correction, facing limitations such as high computational costs, heavy hyperparameter tuning process, and coarse-grained optimization. To address these challenges, we propose a novel two-stage noisy learning framework that enables instance-level optimization through a dynamically weighted loss function, avoiding hyperparameter tuning. To obtain stable and accurate information about noise modeling, we introduce a simple yet effective metric, termed wrong event, which dynamically models the cleanliness and difficulty of individual samples while maintaining computational costs. Our framework first collects wrong event information and builds a strong base model. Then we perform noise-robust training on the base model, using a probabilistic model to handle the wrong event information of samples. Experiments on six synthetic and real-world LNL benchmarks demonstrate our method surpasses state-of-the-art methods in performance, achieves a nearly 75% reduction in storage and computational time, strongly improving model scalability. Our code is available at https://anonymous.4open.science/r/IDO-noisy_label/.
@@ 1 Introduction
To mitigate the negative impact of label noise, many works [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] consider the patterns of the clean and noisy samples shown in the DNN training process, i.e., DNNs first learn clean patterns early in training, then start fitting noise later, and eventually overfit the noisy dataset. However, one key problem of this strategy is that the noisy samples tend to show behavior similar to the clean but difficult samples – the clean examples that are close to the decision boundary of the model. To solve this problem, the most recent approaches [9, 11] try to distinguish noisy examples from hard examples by introducing the hardness of the samples into the loss function. However, these methods have the following limitations.
@@
Figure 2: A comparison between wrong event and loss. The baseline model is ResNet-18 trained on CIFAR-10 with 40% symmetric noise training for 100 epochs. We show the loss distributions (the first row) and wrong event distribution (the second row) during training. The four columns, (a) (b) (c) (d), represent the distributions at epoch 20, 60, 100 and the entire training phase. (a), (b), and (c) represent the early, middle, and late stages, respectively. In (d), the heavy lines represent the median value and the shaded areas are the interquartile ranges. Since wrong events are monotonically increasing based on historical statistics instead of current model prediction, when model overfits the dataset, wrong event values for all samples do not change, rather than converging to zero as loss values typically do. As a result, wrong event can clearly separate noisy data and clean data in all training stages, even if the model fits noise.
@@ 3.2 Observation of wrong event
$\text{wrong event}_{i}=\sum_{t=1}^{T}\mathbb{I}(f_{t}(x_{i})\neq\bar{y}_{i})$ (1)
@@ 3.3 Measuring cleanliness and difficulty
[…]
@@ 4.2 Analysis of loss function
$\mathcal{L_{SIM}}$ is the mean squared error (MSE) between the two predictions. We use $\mathcal{L_{SIM}}$ to increase the loss weight for difficult samples. This encourages the model to focus more on learning from these samples, thereby achieving more robust feature extraction and classification capabilities.
@@ 5 Experiments
Standard ResNet-50 75.5 59.4 41.1 53.4 58.8 68.8 45.9 53.7
SCE ResNet-50 75.2 69.3 55.2 67.0 57.2 69.7 58.6 67.3
TURN ResNet-50 80.7 78.6 73.3 69.4 69.8 70.2 67.5 65.6
ELR ResNet-50 80.9 78.2 73.9 75.6 79.7 70.5 64.2 75.6
CoL ResNet-50 83.1 81.7 79.2 73.6 80.1 70.9 67.3 72.7
DMix ResNet-50 84.3 83.1 80.7 66.6 81.3 73.6 71.4 73.4
UNICON ResNet-50 84.1 83.2 81.1 76.8 82.2 75.6 73.4 74.7
DISC ResNet-50 83.6 82.2 80.1 77.1 82.3 75.7 74.6 75.8
DeFT CLIP-ResNet-50 84.3 82.9 79.9 69.9 82.5 76.8 73.9 74.1
@@
SCE ViT-16/B 91.3 90.4 87.4 73.4 83.6 87.5 86.1 83.2
TURN ViT-16/B 91.0 89.7 88.6 85.1 84.6 86.7 75.3 83.7
ELR ViT-16/B 91.5 90.4 89.7 84.6 91.4 87.2 86.5 87.9
DeFT CLIP-ViT-16/B 92.2 91.2 89.4 72.9 87.6 89.0 72.9 85.2
@@ 5.2 Performance for noisy label learning
Standard ResNet-50 56.8 67.5 69.8 77.8
SCE ResNet-50 61.4 69.6 71.8 78.3
TURN ResNet-50 69.5 68.3 72.9 79.2
DeFT CLIP-ResNet-50 65.5 70.4 73.3 76.6
ELR ResNet-50 69.2 70.5 73.1 80.1
CoL ResNet-50 71.2 72.0 73.0 80.9
DMix ResNet-50 72.9 72.2 74.5 81.7
UNICON ResNet-50 73.1 72.4 74.7 82.5
DISC ResNet-50 73.3 72.1 74.1 82.7
@@
SCE VIT-16/B 76.4 69.5 73.2 83.2
TURN ViT-16/B 78.1 70.3 73.9 83.5
ELR ViT-16/B 79.0 72.2 74.3 83.9
UNICON* CLIP-ViT-16/B 77.7 70.4 - 84.6
LongReMix* CLIP-ViT-16/B 74.0 70.6 - 84.9
ProMix* CLIP-ViT-16/B 76.0 70.7 - 84.5
DeFT* CLIP-ViT-16/B 79.1 72.5 - 85.2
@@ 5.3 Further analysis
Efficiency and scalability of IDO. Existing methods have two drawbacks: high memory demand and low computing efficiency. For instance, DMix and UNICON train two networks simultaneously, which consumes a large amount of memory and limits scalability. Moreover, each sample requires 4 to 8 feed-forward computations, resulting in low efficiency. IDO, in contrast, only requires training a single network using wrong event metric, which can be scaled to larger models. Additionally, each sample only needs two forward computations (two-views), achieving higher efficiency. Fig.4 shows that IDO achieves state-of-the-art performance while achieving high efficiency. Table 4 shows IDO is 4$\times$ faster and 3$\times$ less memory than DMix and UNICON on large models, with better performance on average. For detailed training time analysis, please refer to Appendix D.
@@
Wrong event VS. other metrics. We adjust the BMM fitting distributions of different metrics to compare their performance in noise modeling. We change the duration of stage 1 to evaluate the metrics at different training stages. The results are presented in Table 4. Loss shows optimal performance in the early stage, but the optimal point is hard to identify; Even worse, the performance rapidly degrades in later stages due to noise overfitting. While the EMA-loss (Exponential Moving Average) mitigates noise overfitting in later training phases, it limits the capacity during the initial phase, creating a trade-off between early-stage and late-stage. The visualization about the clean sample selection ability of loss and wrong event is in Appendix C.4. Both forgetting event (FE) and First k-epoch Learning (FkL) show their noise modeling capabilities after the model begins to overfit the noise. However, these metrics need to accumulate over time to show differences, resulting in suboptimal performance in the early stages. Overall, wrong event exhibits superior noise modeling capabilities in all training stages. The experiment results match our theoretical analysis in Section 3.2, confirming the effectiveness of wrong event in identifying noise. More analysis on metrics is in Appendix C.2 C.3.
@@ Appendix B Additional implementation details
Learning Rate $1\times 10^{-2}$ $1\times 10^{-3}$ $1\times 10^{-4}$ $1\times 10^{-2}$ $1\times 10^{-3}$ $1\times 10^{-4}$
Weight Decay $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-4}$
@@ C.1 More empirical observations of wrong event
C.2 Analysis on the variance between wrong event and loss
We theoretically prove that loss variance has no upper bound, but wrong event does, decreasing with training. For epoch $T$, loss $l_{i}^{T}$, $l_{i}^{T+1}\in(0,+\infty)$, wrong event $w_{i}^{T+1}\in[w_{i}^{T},w_{i}^{T}+1]\subseteq[0,T]$, the variation is that
$\frac{w_{i}^{T+1}-w_{i}^{T}}{w_{i}^{T}}\in[0,1/T]\subseteq\frac{l_{i}^{T+1}-l_{i}^{T}}{l_{i}^{T}}\in(-1,+\infty)$
which clearly shows that the variance of wrong event has a upper bound, decreasing over training and the variance of loss has no upper bound, more likely to become large when the model already fits the given label but then predicts another label. We assume that loss variance becomes larger over training, because after the loss converge to loss, i.e., model remembers the sample, forgetting is more likely to bring a mutation to loss.
In experiments, the values of loss and wrong are not in the same range. We normalize metric values and track the range of value changes during training $\delta_{i}(l_{i}^{t})=|l^{t}_{i}-l^{t-1}_{i}|,\delta_{i}(w_{i}^{t})=|w_{i}^{t}-w_{i}^{t-1}|$ and $\text{Max Change}=\mathop{\text{argmax}}\limits_{(x_{i},\bar{y_{i}})\in\mathcal{D}}(\delta_{i}),\text{Avg Change}=\frac{1}{N}\sum_{i=1}^{n}\delta_{i}$. The result is in Figure 7. Loss variance is near the size of the dataset, confirming that loss has no bound. Wrong event variance decreases over training. Besides, due to the mutation of loss, the mean change is larger than wrong event. Experiment results confirm the theoretical analysis regardless of pre-training or not.
@@ C.3 More discussion about different metrics
Researchers have proposed some metrics trying to replace loss because loss is easy to fit noise, such as forgetting event [22], fluctuation event [54], first k-epoch learning [23]. All of them are based on forgetting. In all the following equations, the symbol $\hat{y_{i}}^{t}$ is used to denote the model prediction of sample $(x_{i},\bar{y_{i}})$ at epoch $t$ Forgetting event (FE) is defined as
$\text{forgetting event}_{i}=\sum_{t=1}^{T}(\bar{y_{i}}=\hat{y_{i}}^{t-1})\wedge(\bar{y_{i}}\neq\hat{y_{i}^{t}})$
@@ C.4 Selection ability between wrong event and loss
Table 4 shows that wrong event outperforms loss in test accuracy. To further compare their clean sample selection ability, we calculate F-score along with precision and recall, AUC along with accuracy and precision under pre-training and random initialized settings. Table 9 and Table 10 show the F-score, Table 11 and Table 12 show the AUC value and LABEL:auc-plot shows the AUC-ROC curves during training. All results indicate that wrong event is better than loss in selecting clean samples. Wrong event can clearly model the noise during the whole training phase, while loss only performs well in early stage. In conclusion, wrong event is a simple but effective metric to select clean samples, which can capture noise well even if the model has fitted the noise.

### v4

@@ Abstract
Recent studies indicate that deep neural networks degrade in generalization performance under label noise. Existing methods focus on data selection or label correction, facing limitations such as high computational costs, heavy hyperparameter tuning process, and coarse-grained optimization. To address these challenges, we propose a novel two-stage noisy learning framework that enables instance-level optimization through a dynamically weighted loss function, avoiding hyperparameter tuning. To obtain stable and accurate information about noise modeling, we introduce a simple yet effective metric, termed wrong event, which dynamically models the cleanliness and difficulty of individual samples while maintaining computational costs. Our framework first collects wrong event information and builds a strong base model. Then we perform noise-robust training on the base model, using a probabilistic model to handle the wrong event information of samples. Experiments on six synthetic and real-world LNL benchmarks demonstrate our method surpasses state-of-the-art methods in performance, achieves a nearly 75% reduction in storage and computational time, strongly improving model scalability. Our code is available at https://github.com/iTheresaApocalypse/IDO.
@@ 1 Introduction
To mitigate the negative impact of label noise, many works [6, 7, 8, 9, 10, 11, 12, 13, 14] consider the patterns of the clean and noisy samples shown in the DNN training process, i.e., DNNs first learn clean patterns early in training, then start fitting noise later, and eventually overfit the noisy dataset. However, one key problem of this strategy is that the noisy samples tend to show behavior similar to the clean but difficult samples – the clean examples that are close to the decision boundary of the model. To solve this problem, the most recent approaches [9, 11] try to distinguish noisy examples from hard examples by introducing the hardness of the samples into the loss function. However, these methods have the following limitations.
@@
Figure 2: A comparison between wrong event and loss. The baseline model is ResNet-18 trained on CIFAR-10 with 40% symmetric noise training for 100 epochs. We show the loss distributions (the first row) and wrong event distribution (the second row) during training. The four columns, (a) (b) (c) (d), represent the distributions at epoch 20, 60, 100 and the entire training phase. (a), (b), and (c) represent the early, middle, and late stages, respectively. In (d), the heavy lines represent the mean value and the shaded areas are the interquartile ranges. Since wrong events are monotonically increasing based on historical statistics instead of current model prediction, when model overfits the dataset, wrong event values for all samples do not change, rather than converging to zero as loss values typically do. As a result, wrong event can clearly separate noisy data and clean data in all training stages, even if the model fits noise.
@@ 3.2 Observation of wrong event
$\text{wrong event}_{i}=\sum_{t=1}^{T}\mathbb{I}(\arg\max_{i}f_{t}(x_{i})\neq\bar{y}_{i})$ (1)
@@ 3.3 Measuring cleanliness and difficulty
$\lambda_{k}(w)=F(w;\alpha_{k},\beta_{k})=\frac{\int_{0}^{x}t^{\alpha_{k}-1}(1-t)^{\beta_{k}-1}\,dt}{B(\alpha_{k},\beta_{k})}$ (4)
@@ 4.2 Analysis of loss function
$\mathcal{L_{SIM}}$ is the squared error (SE) between the two predictions. We use $\mathcal{L_{SIM}}$ to increase the loss weight for difficult samples. This encourages the model to focus more on learning from these samples, thereby achieving more robust feature extraction and classification capabilities.
@@ 5 Experiments
Standard RN-50 75.19±0.45 59.41±0.84 41.12±1.35 53.42±1.14 58.84±1.23 68.84±0.51 45.92±1.02 53.72±1.18
SCE RN-50 75.21±0.51 69.37±0.65 55.25±0.71 67.04±0.81 57.18±0.68 69.73±0.41 58.62±0.63 67.36±0.89
TURN RN-50 80.70±0.38 78.68±0.44 73.36±0.53 69.49±0.69 69.82±0.93 70.21±0.59 67.52±0.81 65.61±1.01
ELR RN-50 80.93±0.41 78.22±0.48 73.97±0.59 75.61±0.31 79.77±0.45 70.54±0.51 64.25±0.47 75.62±0.68
CoL RN-50 83.15±0.35 81.76±0.40 79.22±0.59 73.60±0.61 80.18±0.49 70.92±0.63 67.38±0.78 72.71±1.02
DMix RN-50 84.27±0.26 83.08±0.31 80.69±0.36 66.60±0.89 81.37±0.40 73.61±0.36 71.47±0.42 73.39±0.48
UNICON RN-50 84.12±0.29 83.20±0.28 81.12±0.35 76.77±0.39 82.24±0.41 75.61±0.47 73.43±0.56 74.71±0.49
DISC RN-50 83.68±0.36 82.25±0.40 80.12±0.51 77.12±0.44 82.26±0.28 75.73±0.42 74.63±0.57 75.79±0.26
DeFT CLIP-RN-50 84.32±0.21 82.96±0.50 79.92±0.74 69.92±0.81 82.50±0.19 76.81±0.19 73.92±0.37 74.16±0.35
@@
SCE ViT-B 91.26±0.20 90.38±0.29 87.42±0.37 73.46±0.77 83.62±0.44 87.51±0.27 86.12±0.32 83.22±0.51
TURN ViT-B 91.01±0.21 89.75±0.33 88.61±0.35 85.12±0.25 84.67±0.30 86.74±0.31 75.32±0.47 83.72±0.36
ELR ViT-B 91.52±0.16 90.43±0.19 89.74±0.20 84.62±0.23 91.42±0.16 87.22±0.29 86.51±0.34 87.92±0.26
DeFT CLIP-ViT-B 92.17±0.13 91.23±0.17 89.42±0.25 72.96±0.81 87.64±0.29 89.04±0.22 72.93±0.62 85.26±0.31
@@ 5.2 Performance for noisy label learning
Standard ResNet-50 56.71 67.52 69.85 77.84
SCE ResNet-50 61.42 69.65 71.83 78.35
TURN ResNet-50 69.54 68.32 72.98 79.28
DeFT CLIP-ResNet-50 65.53 70.46 73.38 76.62
ELR ResNet-50 69.25 70.52 73.10 80.16
CoL ResNet-50 71.29 72.04 73.09 80.96
DMix ResNet-50 72.91 72.23 74.56 81.72
UNICON ResNet-50 73.12 72.44 74.75 82.48
DISC ResNet-50 73.33 72.15 74.27 82.69
@@
SCE VIT-16/B 76.45 69.52 73.13 83.25
TURN ViT-16/B 78.12 70.31 73.89 83.44
ELR ViT-16/B 79.04 72.23 74.39 83.96
UNICON* CLIP-ViT-16/B 77.76 70.40 - 84.63
LongReMix* CLIP-ViT-16/B 74.09 70.65 - 84.96
ProMix* CLIP-ViT-16/B 76.07 70.79 - 84.55
DeFT* CLIP-ViT-16/B 79.12 72.59 - 85.21
@@ 5.3 Further analysis
[…]
@@
Wrong event VS. other metrics. We adjust the BMM fitting distributions of different metrics to compare their performance in noise modeling. We change the duration of stage 1 to evaluate the metrics at different training stages. The results are presented in Table 4. Loss shows optimal performance in the early stage, but the optimal point is hard to identify; Even worse, the performance rapidly degrades in later stages due to noise overfitting. While the EMA-loss (Exponential Moving Average) mitigates noise overfitting in later training phases, it limits the capacity during the initial phase, creating a trade-off between early-stage and late-stage. The visualization about the clean sample selection ability of loss and wrong event is in Appendix C.4. Both forgetting event (FE) and First k-epoch Learning (FkL) show their noise modeling capabilities after the model begins to overfit the noise. However, these metrics need to accumulate over time to show differences, resulting in suboptimal performance in the early stages. Overall, wrong event exhibits superior noise modeling capabilities in all training stages. The experiment results match our theoretical analysis in Section 3.2, confirming the effectiveness of wrong event in identifying noise. More analysis on metrics is in Appendix C.2 C.3..
@@
Table 6: Comparison of our dynamic weight $\epsilon(w)$ against several fixed-weighting methods on CIFAR-100 with different noise types
Weighting Method Sym. 60% Asym. 40% Inst. 40%
Without $\epsilon=0$ 79.5 70.4 77.3
Fixed $\epsilon=0.25$ 81.3 76.5 83.3
Fixed $\epsilon=0.5$ 80.8 77.5 82.4
Fixed $\epsilon=1$ 80.3 76.8 82.8
Dynamic $\epsilon$ 81.4 78.2 83.8
The behavior of the difficulty-based weighting. IDO relies more heavily on consistency learning under difficult conditions. We conducted new experiments comparing our dynamic weight $\epsilon(w)$ against several fixed-weighting methods (See Table 6). The result supports our core hypothesis, much like the principle behind Focal Loss [33], showing that a weighting term to expand the loss of difficult data is really necessary and better than treating all samples the same. We consider that for easy data, cross-entropy loss serves as a more direct and effective learning signal than consistency loss while consistency loss serves as a effective learning signal for difficult data. Dropping and fixing the difficulty coefficient both decrease the performance. More exploratory experiments are in Appendix E.
6 Limitations and broader impact
6.1 Limitations
6.2 Broader Impact
7 Conclusion
In this work, we introduce IDO, a novel framework that tackles noisy labels through fine-grained, instance-level optimization. Our approach is centered on the wrong event metric, a simple yet highly stable measure of sample cleanliness and difficulty that remains effective throughout training, even post-overfitting. By modeling this metric with a Beta Mixture Model, IDO uses a dynamic loss function that eliminates the need for manual hyperparameter tuning. Experiments confirm that IDO sets a new state-of-the-art in accuracy while being significantly more efficient and scalable than previous methods, offering a practical and robust solution for real-world noisy label learning.
Acknowledgments
Chengliang Chai is supported by National Key Research and Development Program of China (2024YFC3308200), NSFC (62472031), CCF-Baidu Open Fund (CCF-Baidu202402) and Huawei. Ye Yuan is supported by Beijing Natural Science Foundation (L241010), the National Key R&D Program of China (Grant No.2022YFB2702100), and the NSFC (Grant Nos. 61932004, 62225203, U21A20516). Guoren Wang is supported by the NSFC (62427808, U2001211), the Liaoning Revitalization Talents Program (Grant No.XLYC2204005).
@@ Appendix B Additional implementation details
Learning Rate $\times 10^{-2} 1×1$$\times 10^{-3}$ $\times 10^{-4} 1×1$$\times 10^{-2}$ $\times 10^{-3} 1×1$$\times 10^{-4}$
Weight Decay $\times 10^{-5} 1×1$$\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-5}$ $\times 10^{-5} 1×1$$\times 10^{-4}$
@@ C.1 More empirical observations of wrong event
C.2 Analysis on the change rate between wrong event and loss
We provide an analysis about the change rate between wrong event and loss. The loss change rate is $\frac{\delta_{i}^{T+1}-\delta_{i}^{T}}{\delta_{i}^{T}}\in[-\infty,+\infty]$ in every epoch $T$ with unbounded limits and frequently unpredictable mutations. A usually seen phenomenon named forgetting event [18] (last epoch right, this epoch wrong) can cause the probability of given label to plummet from a high probability (e.g. $\sim 0.9$) to a small probability (e.g. $\sim 0.1$), causing the loss to skyrocket and the relative change to become enormous. This can happen randomly and repeatedly for all samples.
However, for wrong event, the change rate needs more discussion. Firstly, we have $w_{i}^{T+1}=w_{i}^{T}+\delta_{i}^{T+1},\delta_{i}^{T+1}\in\{0,1\}$ and $w_{i}^{T}\in[0,\ldots,T]$, meaning the wrong event value is monotonically increasing. The change rate is $\frac{\delta_{i}^{T+1}}{w_{i}^{T}}$. We analyze this in two distinct cases:
The “0 to 1” Transition: When a sample is misclassified for the very first time, $w_{i}^{T}$ transitions from 0 to 1. At this point, the relative change $\frac{1}{0}$ is not defined. This is a singular, one-time event for any given sample, marking its initial entry into the set of “ever-misclassified” samples.
All Subsequent Changes ($w_{i}^{T}\geq 1$): For every subsequent misclassification, the denominator $w_{i}^{T}$ is at least 1. The relative change is $\frac{\delta_{i}^{T+1}}{w_{i}^{T}}$. This value is strictly bounded within the range [0,1] far smaller than the change rate of loss. More importantly, as the model continues to misclassify a sample, $w_{i}$ increases, causing the maximum possible relative change ($\frac{1}{w_{i}^{T}}$) to decrease. The metric is therefore self-stabilizing.
The relative change of wrong event is well-behaved, bounded, and self-stabilizing after a single initial event. Conversely, the relative change of loss is unbounded and perpetually susceptible to explosive volatility. This fundamental difference in their rate of change is why wrong event serves as a far more reliable and stable signal for modeling sample characteristics in noisy environments.
This analysis provides a strong theoretical basis for the empirical stability we observed in Figure 7. In experiments, the values of loss and wrong are not in the same range. We normalize metric values and track the range of value changes during training $\delta_{i}(l_{i}^{t})=|l^{t}_{i}-l^{t-1}_{i}|,\delta_{i}(w_{i}^{t})=|w_{i}^{t}-w_{i}^{t-1}|$ and $\text{Max Change}=\mathop{\text{argmax}}\limits_{(x_{i},\bar{y_{i}})\in\mathcal{D}}(\delta_{i}),\text{Avg Change}=\frac{1}{N}\sum_{i=1}^{n}\delta_{i}$. The result is in Figure 7. Loss variance is near the size of the dataset, confirming that loss has no bound. Wrong event variance decreases over training. Besides, due to the mutation of loss, the mean change is larger than wrong event. Experiment results confirm the theoretical analysis regardless of pre-training or not.
@@ C.3 More discussion about different metrics
Researchers have proposed some metrics trying to replace loss because loss is easy to fit noise, such as forgetting event [18], fluctuation event [51], first k-epoch learning [19]. All of them are based on forgetting. In all the following equations, the symbol $\hat{y_{i}}^{t}$ is used to denote the model prediction of sample $(x_{i},\bar{y_{i}})$ at epoch $t$ . Forgetting event (FE) is defined as
[…]
@@
Figure 8: The AUC-ROC curves of loss and wrong event. The experiment is conducted with random initialized ResNet-18 on CIFAR-10 (left) and pre-trained ResNet-50 on CIFAR-100 (right) under Inst.40% noise. The curves show the dynamic selection curves over epochs. We visualize the ability of clean sample selection in early, midterm and later training stages. The curves show that wrong event has a larger AUC value than loss and show good performance during the whole training phase, which represents a better ability of clean sample selection.
@@ C.4 Selection ability between wrong event and loss
Table 4 shows that wrong event outperforms loss in test accuracy. To further compare their clean sample selection ability, we calculate F-score along with precision and recall, AUC along with accuracy and precision under pre-training and random initialized settings. Table 10 and Table 11 show the F-score, Table 12 and Table 13 show the AUC value and Figure 8 shows the AUC-ROC curves during training. All results indicate that wrong event is better than loss in selecting clean samples. Wrong event can clearly model the noise during the whole training phase, while loss only performs well in early stage. In conclusion, wrong event is a simple but effective metric to select clean samples, which can capture noise well even if the model has fitted the noise.